A particle of mass m charge q acceleration a then write it's intensity.
1
Expert's answer
2020-09-30T12:28:06-0400

Solution

When a particle of mass m and charge q acceleration a emits radiation with intensity

"I=(constant). \frac{q^2 a^2 \sin^2\theta }{r^2}"